# A Large Recurrent Action Model: xLSTM enables Fast Inference for Robotics Tasks

### v1

@@ 7 Reproducibility
Appendix A Environments & Datasets
A.1 General
@@
A.2 Atari
The Arcade Learning Environment (ALE) [Bellemare et al.,, 2013] is the standard benchmark for evaluating RL agents and consists of 57 Atari games. Input observations in Atari are RGB images, but as is standard practice we gray-scale and crop frames ($|\mathcal{S}|=1\times 64\times 64$). There are 18 discrete action across all 57 Atari games ($|\mathcal{A}|=18$), but individual games may use only use a subset of these actions. Furthermore, we adopt the standard Atari recipe as used in prior works, including a frame skip of 4, maximum number of no-ops of 30, resetting on life loss, and reward clipping to $[-1,1]$ [Mnih et al.,, 2015; Hessel et al.,, 2017].
Tasks. Similar to Lee et al., [2022], we assign 41 games to the training set, and 5 additional tasks to the hold-out set. The 41 training tasks include:
@@ A.2 Atari
Dataset. For Atari, we leverage the DQN-Replay dataset released by Agarwal et al., [2020]. The dataset contains the trajectories seen over the entire training of the DQN agent (50M frames), We extract a subset of the last 5M transitions for every task, amounting to 205M transitions in total for the 41 training tasks. The number of episodes, the episodes lengths and total achieved rewards vary across tasks, as shown in Table 2.
A.3 Meta-World
The Meta-World benchmark [Yu et al., 2020a, ] consists of 50 manipulations tasks using a Sawyer robotic arm, ranging from opening or closing windows, to pressing buttons. Meta-World is based on the MuJoCo physics engine [Todorov et al., 2012a, ]. Observations in Meta-World are 39-dimensional continuous vectors ($|\mathcal{S}|=1\times 64\times 39$), and actions are represented by 6 continuous dimensions ($|\mathcal{A}|=18$) in range $[-1,1]$. All tasks share a common action and state space. Following Wolczyk et al., [2021] and Schmied et al., 2024a , we limit the episode lengths to 200 interactions.
Tasks. We follow Yu et al., 2020a and split the 50 Meta-World tasks into 45 training tasks (MT45) and 5 evaluation tasks (MT5).
@@ A.3 Meta-World
Dataset. For Meta-World, we use the datasets released by [Schmied et al., 2024a, ], which contain 2M transitions per tasks and consequently 90M transitions in total for the training set. All episodes last for 200 environment interaction steps, and consequently there are 10K episodes for every task. For detailed dataset statistics per task, we refer to their publication.
@@
Figure 9: Illustration of the four supported robot arms in Composuite [Mendez et al.,, 2022].
A.4 DMControl
The DMControl benchmark [Tassa et al.,, 2018] consists of 30 different robotic tasks. Unlike Meta-World, the benchmark contains robots with different morphologies instead of a single common Sawyer arm. Due to the different robot morphologies, the state, and action spaces vary across tasks ($3\leq|\mathcal{S}|\leq 24$, $1\leq|\mathcal{A}|\leq 6$), with all actions in range $[-1,1]$.
Tasks. We do not use all 30 tasks contained in the DMControl benchmark, but select 16 of the 30 tasks that have been used in prior works [Hafner et al.,, 2019; Schmied et al., 2024a, ; Schmied et al., 2024b, ], which we refer to as DMC11 and DMC5 respectively.
@@ A.4 DMControl
Dataset. For DMControl, we generate 10M transitions per task by training task-specific SAC [Haarnoja et al.,, 2018] agents, using the same setup as Schmied et al., 2024a . Episodes in all DMControl tasks last for 1000 environment steps and per time-step a maximum reward of +1 can be achieved, which results in a maximum reward of 1000 per episode. Consequently, our training set contains 10K episodes per tasks, amounting to 110K episodes and 110M transitions in total across all tasks. We list the dataset statistics for all 11 tasks in Table 3.
@@
A.5 Composuite
The Composuite benchmark [Mendez et al.,, 2022], is a robotics benchmark for grasping and object manipulation. The benchmark is implemented on top of robotsuite [Zhu et al.,, 2020], which in turn leverages the MuJoCo simulator under the hood [Todorov et al., 2012b, ]. Composuite contains a mix of 4 simulated robot arms: IIWA, Jaco, Gen3, and Panda (see Figure 9). All arms share a common state and action space containing 93 continuous state dimensions and 8 continuous action dimensions, respectively ($|\mathcal{S}|=93$, $|\mathcal{A}|=8$).
Tasks. CompoSuite is designed as a compositional multi-task benchmark for RL, in which a particular robot manipulates a particular object given an objective, while avoiding obstacles. Overall, there are 4 robots arms, 4 objects, 4 obstacles, and 4 task objectives. This results in 256 possible robot/object/objective/obstacles combinations. For our experiments, we assign 240 tasks to the training set and use the remaining 16 tasks as hold-out set (Panda and Object_Wall) combinations. For a list of all 256 tasks, we refer to Mendez et al., [2022].
Dataset. For Composuite, we leverage the datasets released by Hussing et al., [2023]. For every task, we select 2000 episodes, which last on average for 500 steps. This amounts to 1M transitions per task, and 240M transitions across all 240 training tasks. For dataset statistics, we refer to Hussing et al., [2023].
A.6 Mimicgen
Similar to Composuite, Mimicgen [Mandlekar et al.,, 2023] is based on robosuite and the MuJoCo simulator. Mimicgen is designed for automatically synthesizing large-scale datasets from only a handful of human demonstrations. Observations in Mimicgen can be represented as images (from multiple cameras) or low dimensional continuous states. For our experiments, we opt for the low-dimensional state representation to simplify learning. Therefore, observations and actions are represented by 37-dimensional and 7-dimensional continuous vectors, respectively ($|\mathcal{S}|=37$, $|\mathcal{A}|=7$). Similar to Composuite, Mimicgen supports 4 different robot arms: Panda, IIWA, Sawyer, and UR5e (see Figure 10).
@@ A.6 Mimicgen
Figure 10: Illustration of the four supported robot arms in Mimicgen [Mandlekar et al.,, 2023] solving the stack-three task.
@@
Dataset. Mandlekar et al., [2023] released dataset for the 24 tasks using the default robot arm Panda. To increase the dataset diversity, we additionally generated data for the remaining 3 robot arms. However, not all data generation runs produce successful trajectories, and we discard with too few successful trajectories. Our final dataset for Mimicgen contains 83 training and 2 evaluation tasks. For each task, we collect 1000 successful demonstrations (we do not include unsuccessful trajectories). Episode lengths vary across tasks, ranging from 260 to 850 environment steps.
A.7 Procgen
Procgen benchmark consists of 16 procedurally-generated video games [Cobbe et al., 2020b, ]. Observations in Procgen are RGB images of dimension $3\times 64\times 64$. However, for training efficiency, we apply gray-scaling to image observations ($|\mathcal{S}|=1\times 64\times 64$). All 16 environments share a common action space of 15 discrete actions ($|\mathcal{A}|=16$). Procgen is designed to test the generalization abilities of RL agents. Consequently, procedural generation is employed to randomize background and colors, while retaining the game dynamics.
Tasks. Following prior works [Raparthy et al.,, 2023; Schmied et al., 2024b, ], we assign 12 and 4 tasks to training and hold-out set, respectively. The 12 training tasks are:
@@ A.7 Procgen
Dataset. We leverage the datasets released by Schmied et al., 2024b , which contain 20M transitions per task. The datasets were generated by recording all transitions observed by training RL agents for 25M steps, followed by uniform subsampling to 20M transitions. Consequently, the dataset contains mixed quality trajectories ranging from random (beginning of training) to expert (end of training). We list the dataset statistics for all 16 tasks in Table 4.
@@
Appendix B Experimental & Implementation Details
B.1 Training & Evaluation.
In our experiments, we compare two variants of xLSTM, Mamba and DT. For our main experiments in Section 4.2, we train all models for 200K updates, and evaluate after every 50K update steps. We report the mean and 95% confidence intervals over three seeds in our experiments, as suggested by Agarwal et al., [2021]. For every evaluation tasks, we take the average of 3 evaluation seeds.
We train our agents with a batch size of 128 and gradient accumulation across the 6 domains, such that every domain is represented with the same proportion. Consequently, the effective batch size is 768. We use a learning rate of $1e^{-4}$, 4000 linear warm-up steps followed by a cosine decay to $1e^{-6}$, and train using the AdamW optimizer [Loshchilov and Hutter,, 2018]. In addition, we employ gradient clipping of 0.25, weight decay of 0.01 for all models. We do not employ Dropout, as is standard practice in DTs, as we found that it negatively affects performance (see Section 4.3). We use separate reward scales of 200, 100 and 20 for Meta-World, DMControl and Atari, respectively. Furthermore, for all domains, we set the target return to the maximum return achieved for a particular task in the training datasets. This is particularly useful for domains, where the maximum returns differ heavily across tasks (e.g., Atari). We list all hyperparameters in Table 5.
Table 5: Hyperparameters for RA-DT.
@@ Appendix B Experimental & Implementation Details
By default, we train all models with a context length $C=50$ timesteps. For every timestep there are three tokens (s/rt/r) and consequently, the effective context length is 150. We found that performance improves for longer context lengths (see Section D.1), but limit our experiments to $C=50$ to reduce the computational cost.
B.3 Model Architectures.
@@
We train models across 4 models sizes: 16M, 48M, 110M, and 206M. We follow Lee et al., [2022] in selecting the number of layers and hidden dimensions. For xLSTM and Mamba, we use twice the number of layers blocks to match the number of parameters of the Transformer [Beck et al.,, 2024; Gu et al.,, 2024] (see Table 6) For our xLSTM [7:1] variant, which contains sLSTM blocks, we strive to maintain the same ratio as proposed by Beck et al., [2024]. Not all our model sizes are divisible by 8 and only the 16M and 110M models exhibit the exact 7:1 ratio of mLSTM to sLSTM blocks. For consistency, however, we maintain the same notation as Beck et al., [2024]. We place sLSTM blocks at positions [1], [1, 3], [1, 3], and [1, 3, 5] for the 16M, 48M, 110M, 206M, respectively.
@@
Across backbones, we use linear layers to encode continuous states, reward returns-to-go, similar to Chen et al., [2021]. The maximal state-dimension across continuous control environments is 204 in our experiments. To use a shared linear embedding layer for continuous states, we pad states that have lower number of dimensions to 204 dimensions using zeros. To encode image inputs on visual domains, we use the IMPALA-CNN proposed by Espeholt et al., [2018] and adopted by previous works on Procgen [Cobbe et al., 2020b, ] and Atari [Schmidt and Schmied,, 2021; Schwarzer et al.,, 2023]. Consequently, we do not make use of discretization of continuous states or patchification of images. This design choice significantly reduces the sequence length to only three tokens per time-step (see Appendix B.2) and consequently results in faster inference.
@@
For continuous actions, we make use of discretization and discretize of every action dimension into 256 uniformly-spaced bins, similar to Reed et al., [2022] and Brohan et al., 2023b . We experimented with lower/higher number of bins, but did not observe a benefit beyond 256 bins. Consequently, this resolution is sufficient for the environments we consider. We use a shared action head to predict the action bins of all continuous dimensions jointly. The maximum number of continuous action dimensions is 8 in our experiments and consequently the number of discrete action classes is 2048. In addition, there are 18 discrete actions originating from Atari and Procgen. Therefore, our action head learns to predict the correct action among the 2066 discrete classes. We opt for the shared action head representation, as this further speeds up inference and does not require autoregressive action prediction.
@@
For the Transformer baseline, we use global positional embeddings similar to Chen et al., [2021]. For the recurrent backbones, we do not make use of positional encodings.
B.4 Hardware & Training Times.
@@
We train all our models on a server equipped with 4 A100 GPUs. We use distributed data parallel to distribute the workload, as supported in PyTorch [Paszke et al.,, 2019]. Training times range from 5 hours for the smallest DT model to 30 hours for the largest Mamba model. Throughout all our experiments, we use mixed precision training [Micikevicius et al.,, 2017] as supported in PyTorch to speed up training time.
@@
Appendix C Additional Results
C.1 Training Tasks
In Figures 11 and 12, we report the normalized scores obtained per domain and the average learning curves across tasks for all four model sizes.

### v2

@@ Acknowledgements
Appendix A Reproducibility Statement
We make the code-base used for our experiments publicly available, and release the datasets we generated. Both are available at: https://github.com/ml-jku/LRAM. As part of this submission, we also include the source code in the supplementary material. We describe the environments we use for our experiments and provide dataset statistics in Appendix B. Furthermore, in Appendix C, we provide implementation details for all methods and a list of hyperparameters used for our experiments. In Appendix D, we present additional figures that accompany our results in the main text (e.g., all model sizes). Finally, in Appendices E and F, we provide further details on the conducted ablation studies and the embedding space analysis, respectively.
Appendix B Environments & Datasets
B.1 General
@@
B.2 Atari
The Arcade Learning Environment (ALE) (Bellemare et al., 2013) is the standard benchmark for evaluating RL agents and consists of 57 Atari games. Input observations in Atari are RGB images, but as is standard practice we gray-scale and crop frames ($|\mathcal{S}|=1\times 64\times 64$). There are 18 discrete action across all 57 Atari games ($|\mathcal{A}|=18$), but individual games may use only use a subset of these actions. Furthermore, we adopt the standard Atari recipe as used in prior works, including a frame skip of 4, maximum number of no-ops of 30, resetting on life loss, and reward clipping to $[-1,1]$ (Mnih et al., 2015; Hessel et al., 2017).
Tasks. Similar to Lee et al. (2022), we assign 41 games to the training set, and 5 additional tasks to the hold-out set. The 41 training tasks include:
@@ B.2 Atari
Dataset. For Atari, we leverage the DQN-Replay dataset released by Agarwal et al. (2020). The dataset contains the trajectories seen over the entire training of the DQN agent (50M frames), We extract a subset of the last 5M transitions for every task, amounting to 205M transitions in total for the 41 training tasks. The number of episodes, the episodes lengths and total achieved rewards vary across tasks, as shown in Table 2.
B.3 Meta-World
The Meta-World benchmark (Yu et al., 2020a) consists of 50 manipulations tasks using a Sawyer robotic arm, ranging from opening or closing windows, to pressing buttons. Meta-World is based on the MuJoCo physics engine (Todorov et al., 2012a). Observations in Meta-World are 39-dimensional continuous vectors ($|\mathcal{S}|=1\times 64\times 39$), and actions are represented by 6 continuous dimensions ($|\mathcal{A}|=18$) in range $[-1,1]$. All tasks share a common action and state space. Following Wolczyk et al. (2021) and Schmied et al. (2024a), we limit the episode lengths to 200 interactions.
Tasks. We follow Yu et al. (2020a) and split the 50 Meta-World tasks into 45 training tasks (MT45) and 5 evaluation tasks (MT5).
@@ B.3 Meta-World
Dataset. For Meta-World, we use the datasets released by (Schmied et al., 2024a), which contain 2M transitions per tasks and consequently 90M transitions in total for the training set. All episodes last for 200 environment interaction steps, and consequently there are 10K episodes for every task. For detailed dataset statistics per task, we refer to their publication.
@@
Figure 8: Illustration of the four supported robot arms in Composuite (Mendez et al., 2022).
B.4 DMControl
The DMControl benchmark (Tassa et al., 2018) consists of 30 different robotic tasks. Unlike Meta-World, the benchmark contains robots with different morphologies instead of a single common Sawyer arm. Due to the different robot morphologies, the state, and action spaces vary across tasks ($3\leq|\mathcal{S}|\leq 24$, $1\leq|\mathcal{A}|\leq 6$), with all actions in range $[-1,1]$.
Tasks. We do not use all 30 tasks contained in the DMControl benchmark, but select 16 of the 30 tasks that have been used in prior works (Hafner et al., 2019; Schmied et al., 2024a, b), which we refer to as DMC11 and DMC5 respectively.
@@ B.4 DMControl
Dataset. For DMControl, we generate 10M transitions per task by training task-specific SAC (Haarnoja et al., 2018) agents, using the same setup as Schmied et al. (2024a). Episodes in all DMControl tasks last for 1000 environment steps and per time-step a maximum reward of +1 can be achieved, which results in a maximum reward of 1000 per episode. Consequently, our training set contains 10K episodes per tasks, amounting to 110K episodes and 110M transitions in total across all tasks. We list the dataset statistics for all 11 tasks in Table 3.
@@
B.5 Composuite
The Composuite benchmark (Mendez et al., 2022), is a robotics benchmark for grasping and object manipulation. The benchmark is implemented on top of robotsuite (Zhu et al., 2020), which in turn leverages the MuJoCo simulator under the hood (Todorov et al., 2012b). Composuite contains a mix of 4 simulated robot arms: IIWA, Jaco, Gen3, and Panda (see Figure 8). All arms share a common state and action space containing 93 continuous state dimensions and 8 continuous action dimensions, respectively ($|\mathcal{S}|=93$, $|\mathcal{A}|=8$).
Tasks. CompoSuite is designed as a compositional multi-task benchmark for RL, in which a particular robot manipulates a particular object given an objective, while avoiding obstacles. Overall, there are 4 robots arms, 4 objects, 4 obstacles, and 4 task objectives. This results in 256 possible robot/object/objective/obstacles combinations. For our experiments, we assign 240 tasks to the training set and use the remaining 16 tasks as hold-out set (Panda and Object_Wall) combinations. For a list of all 256 tasks, we refer to Mendez et al. (2022).
Dataset. For Composuite, we leverage the datasets released by Hussing et al. (2023). For every task, we select 2000 episodes, which last on average for 500 steps. This amounts to 1M transitions per task, and 240M transitions across all 240 training tasks. For dataset statistics, we refer to Hussing et al. (2023).
B.6 Mimicgen
Similar to Composuite, Mimicgen (Mandlekar et al., 2023) is based on robosuite and the MuJoCo simulator. Mimicgen is designed for automatically synthesizing large-scale datasets from only a handful of human demonstrations. Observations in Mimicgen can be represented as images (from multiple cameras) or low dimensional continuous states. For our experiments, we opt for the low-dimensional state representation to simplify learning. Therefore, observations and actions are represented by 37-dimensional and 7-dimensional continuous vectors, respectively ($|\mathcal{S}|=37$, $|\mathcal{A}|=7$). Similar to Composuite, Mimicgen supports 4 different robot arms: Panda, IIWA, Sawyer, and UR5e (see Figure 9).
@@ B.6 Mimicgen
Figure 9: Illustration of the four supported robot arms in Mimicgen (Mandlekar et al., 2023) solving the stack-three task.
@@
Dataset. Mandlekar et al. (2023) released dataset for the 24 tasks using the default robot arm Panda. To increase the dataset diversity, we additionally generated data for the remaining 3 robot arms. However, not all data generation runs produce successful trajectories, and we discard with too few successful trajectories. Our final dataset for Mimicgen contains 83 training and 2 evaluation tasks. For each task, we collect 1000 successful demonstrations (we do not include unsuccessful trajectories). Episode lengths vary across tasks, ranging from 260 to 850 environment steps.
B.7 Procgen
Procgen benchmark consists of 16 procedurally-generated video games (Cobbe et al., 2020a). Observations in Procgen are RGB images of dimension $3\times 64\times 64$. However, for training efficiency, we apply gray-scaling to image observations ($|\mathcal{S}|=1\times 64\times 64$). All 16 environments share a common action space of 15 discrete actions ($|\mathcal{A}|=16$). Procgen is designed to test the generalization abilities of RL agents. Consequently, procedural generation is employed to randomize background and colors, while retaining the game dynamics.
Tasks. Following prior works (Raparthy et al., 2023; Schmied et al., 2024b), we assign 12 and 4 tasks to training and hold-out set, respectively. The 12 training tasks are:
@@ B.7 Procgen
Dataset. We leverage the datasets released by (Schmied et al., 2024b), which contain 20M transitions per task. The datasets were generated by recording all transitions observed by training RL agents for 25M steps, followed by uniform subsampling to 20M transitions. Consequently, the dataset contains mixed quality trajectories ranging from random (beginning of training) to expert (end of training). We list the dataset statistics for all 16 tasks in Table 4.
@@
Appendix C Experimental & Implementation Details
C.1 Training & Evaluation.
In our experiments, we compare two variants of xLSTM, Mamba and DT. For our main experiments in Section 4.2, we train all models for 200K updates, and evaluate after every 50K update steps. We report the mean and 95% confidence intervals over three seeds in our experiments, as suggested by Agarwal et al. (2021). For every evaluation tasks, we take the average of 3 evaluation seeds.
We train our agents with a batch size of 128 and gradient accumulation across the 6 domains, such that every domain is represented with the same proportion. Consequently, the effective batch size is 768. We use a learning rate of $1e^{-4}$, 4000 linear warm-up steps followed by a cosine decay to $1e^{-6}$, and train using the AdamW optimizer (Loshchilov & Hutter, 2018). In addition, we employ gradient clipping of 0.25, weight decay of 0.01 for all models. We do not employ Dropout, as is standard practice in DTs, as we found that it negatively affects performance (see Section 4.3). We use separate reward scales of 200, 100 and 20 for Meta-World, DMControl and Atari, respectively. Furthermore, for all domains, we set the target return to the maximum return achieved for a particular task in the training datasets. This is particularly useful for domains, where the maximum returns differ heavily across tasks (e.g., Atari). We list all hyperparameters in Table 5.
Table 5: Hyperparameters for LRAM.
@@ Appendix C Experimental & Implementation Details
By default, we train all models with a context length $C=50$ timesteps. For every timestep there are three tokens (s/rt/r) and consequently, the effective context length is 150. We found that performance improves for longer context lengths (see Section E.1), but limit our experiments to $C=50$ to reduce the computational cost.
C.3 Model Architectures.
@@
We train models across 4 models sizes: 16M, 48M, 110M, and 206M. We follow Lee et al. (2022) in selecting the number of layers and hidden dimensions. For xLSTM and Mamba, we use twice the number of layers blocks to match the number of parameters of the Transformer (Beck et al., 2024; Gu et al., 2024) (see Table 6) For our xLSTM [7:1] variant, which contains sLSTM blocks, we strive to maintain the same ratio as proposed by Beck et al. (2024). Not all our model sizes are divisible by 8 and only the 16M and 110M models exhibit the exact 7:1 ratio of mLSTM to sLSTM blocks. For consistency, however, we maintain the same notation as (Beck et al., 2024). We place sLSTM blocks at positions [1], [1, 3], [1, 3], and [1, 3, 5] for the 16M, 48M, 110M, 206M, respectively.
@@
Across backbones, we use linear layers to encode continuous states, reward returns-to-go, similar to Chen et al. (2021). The maximal state-dimension across continuous control environments is 204 in our experiments. To use a shared linear embedding layer for continuous states, we pad states that have lower number of dimensions to 204 dimensions using zeros. To encode image inputs on visual domains, we use the IMPALA-CNN proposed by Espeholt et al. (2018) and adopted by previous works on Procgen (Cobbe et al., 2020a) and Atari (Schmidt & Schmied, 2021; Schwarzer et al., 2023). Consequently, we do not make use of discretization of continuous states or patchification of images. This design choice significantly reduces the sequence length to only three tokens per time-step (see Appendix C.2) and consequently results in faster inference.
@@
For continuous actions, we make use of discretization and discretize of every action dimension into 256 uniformly-spaced bins, similar to Reed et al. (2022) and Brohan et al. (2023b). We experimented with lower/higher number of bins, but did not observe a benefit beyond 256 bins. Consequently, this resolution is sufficient for the environments we consider. We use a shared action head to predict the action bins of all continuous dimensions jointly. The maximum number of continuous action dimensions is 8 in our experiments and consequently the number of discrete action classes is 2048. In addition, there are 18 discrete actions originating from Atari and Procgen. Therefore, our action head learns to predict the correct action among the 2066 discrete classes. While different environments may have different action dimensions, the model predicts all action dimensions jointly. At inference time, the number of action dimensions of the current environment is known, and we extract the respective dimensions from the joint predictions. We opt for the shared action head representation, as this further speeds up inference and does not require autoregressive action prediction.
@@
For the Transformer baseline, we use global positional embeddings similar to Chen et al. (2021). For the recurrent backbones, we do not make use of positional encodings.
C.4 Hardware & Training Times.
@@
We train all our models on a server equipped with 4 A100 GPUs. We use distributed data parallel to distribute the workload, as supported in PyTorch (Paszke et al., 2019). Training times range from 5 hours for the smallest DT model to 30 hours for the largest Mamba model. Throughout all our experiments, we use mixed precision training (Micikevicius et al., 2017) as supported in PyTorch to speed up training time.
@@
Appendix D Additional Results
D.1 Training Tasks
In Figures 10 and 11, we report the normalized scores obtained per domain and the average learning curves across tasks for all four model sizes.